OpenStudy (anonymous):

Factoring...help?~Will reward 4a^2=9/64 x^n+6 +x^n+2 +x^n

zepdrix (zepdrix):

\[\large 4a^2=\frac{9}{64}\]Subtracting 9/64 from each side gives us,\[\large 4a^2-\frac{9}{64}=0\] 9/64 is actually a perfect square divided by a perfect square, let's write it like that. And we'll do the same with the 4.\[\large 2^2a^2-\frac{3^2}{8^2}\]Which we can write as,\[\large (2a)^2-\left(\frac{3}{8}\right)^2\] We haven't factored yet. That part won't be too bad. But so far, do you understand what I did? I rewrote some stuff as squares so we can get through the next part easily.

OpenStudy (anonymous):

Yes

zepdrix (zepdrix):

What we have here is the DIFFERENCE OF SQUARES. There is a handy dandy formula for factoring the difference of squares.\[\large x^2-y^2=(x-y)(x+y)\] In our case, \(2a\) is our X that's being squared, while \(\huge \frac{3}{8}\) is our Y. \[\large (2a)^2-\left(\frac{3}{8}\right)^2 \qquad \rightarrow \qquad \left(2a-\frac{3}{8}\right)\left(2a+\frac{3}{8}\right)\]

zepdrix (zepdrix):

Are you suppose to solve for \(a\)? Or did they just want it factored?

OpenStudy (anonymous):

It said to solve each equation by factoring and the answers they had were -3/16 and 3/16

OpenStudy (anonymous):

I was wondering how they got there...

zepdrix (zepdrix):

From here,\[\large \left(2a-\frac{3}{8}\right)\left(2a+\frac{3}{8}\right)=0\]We can apply the Zero Factor Property. We'll set each individual factor equal to 0 and solve for \(a\).\[\large 2a-\frac{3}{8}=0, \qquad \qquad 2a+\frac{3}{8}=0\]

zepdrix (zepdrix):

Looking at the first one, we'll add 3/8 to each side.\[\large 2a=\frac{3}{8}\]We'll divide both sides by 2, giving us,\[\large a=\frac{3}{16}\]

zepdrix (zepdrix):

We can do the same with the other factor. If you're having trouble with the factoring, you can actually solve this problem without factoring. As a first step, take the square root of both sides.\[\large \large 4a^2=\frac{9}{64} \qquad \rightarrow \qquad 2a= \pm \frac{3}{8}\]Then divide both sides by 2,\[\large a=\pm \frac{3}{16}\]

OpenStudy (anonymous):

^-^ I get it now ;-;

zepdrix (zepdrix):

That certainly isn't the way they wanted you to do it, but it's much easier :D heh